The Shoe Outlet has paid annual dividends of $0.80, $0.85, $1.05, $1.25, and $1.50 per share

over the last five years, respectively. The stock is currently selling for $56.23 a share. What is

the cost of equity?
Answer #1

Growth rate 1=(.85-.80)/.80=6.25%

Growth rate 1=(1.05-.85)/.85=23.53%

Growth rate 1=(1.25-1.05)/1.05=19.05%

Growth rate 1=(1.50-1.25)/1.25=20.00%

Growth Rate =(6.25%+23.53%+19.05%+20%)/4=17.21%

cost of equity =D1/Selling Price+growth

=(1.50*(1+17.21%))/56.23+17.21%

cost of equity=20.34%